A spacecraft with a proper length of 300 m takes 0.75 ms to pass an Earth-based observer. Determine the speed of the spacecraft as measured by the Earth observer.


1
Expert's answer
2021-11-14T17:09:16-0500

"l_0=300 \ (m)"

"t=0.75\cdot10^{-6}\ (s)"



"l=l_0\sqrt{1-v^2/c^2}\to l^2=l_0^2(1-v^2/c^2)"


"l=vt\to l^2=\frac{v^2}{c^2}(ct)^2" . After equating these two expressions we will get


"l_0^2(1-v^2/c^2)=\frac{v^2}{c^2}(ct)^2\to l_0^2=\frac{v^2}{c^2}(l_0^2+(ct)^2)"


"300^2=\frac{v^2}{c^2}(300^2+(c\cdot0.00000075)^2)\to v=0.8c" . Answer